OpenStudy (pigsruleyeah):

The following data shows the weight, in pounds, of 6 bags: WILL FAN AND MEDAL 6, 4, 8, 7, 8, 9 What is the value of the mean absolute deviation of the weight of the bags, and what does it represent about the weight of a bag? 1.8 pounds; on average, weight of a bag varies 1.8 pounds from the mean of 7 pounds 1.3 pounds; on average, weight of a bag varies 1.3 pounds from the mean of 7 pounds 1.3 pounds; the weight of 50% of the bags is greater than 1.3 pounds 1.8 pounds; the weight of 50% of the bags is greater than 1.8 pounds

OpenStudy (pigsruleyeah):

HELP

OpenStudy (pigsruleyeah):

helppp meee

OpenStudy (pigsruleyeah):

@PokemonFanatic40

OpenStudy (pigsruleyeah):

@PokemonFanatic40

OpenStudy (pokemonfanatic40):

what?

OpenStudy (pigsruleyeah):

helpppppppp

OpenStudy (pokemonfanatic40):

i am hang on

OpenStudy (pigsruleyeah):

HELPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP

OpenStudy (pokemonfanatic40):

it is 1.3

OpenStudy (pigsruleyeah):

ikt but thee restt

OpenStudy (pokemonfanatic40):

i don't know ok

OpenStudy (pigsruleyeah):

ok @Mystiz

OpenStudy (pigsruleyeah):

help

OpenStudy (mystiz):

ok not to hard at all

OpenStudy (pigsruleyeah):

i know its 1.3 i dont know the rest

OpenStudy (mystiz):

6+ 4+ 8+ 7+ 8+ 9=42 ÷ 6= 7

OpenStudy (mystiz):

7 is the average

OpenStudy (pigsruleyeah):

k

OpenStudy (mystiz):

or mean

OpenStudy (pigsruleyeah):

so its b

OpenStudy (mystiz):

could be i am not the math wiz but my sister is

OpenStudy (pigsruleyeah):

ask her

OpenStudy (mystiz):

she will solve this thing in less then 60 seconds with everything explained

OpenStudy (mystiz):

she is not here